# MaGRITTe: Manipulative and Generative 3D Realization from Image, Topview and Text

### v1

@@ 2 Related Work
3D scene generation involves the creation of a model of a 3D space that includes objects and backgrounds, based on user-specified conditions. In recent years, the use of generative models, such as VAE [30, 44], GAN [19], autoregressive models [57], and diffusion models [50, 25], has made rapid progress. There are methods to generate a 3D scene from random variables [37, 8], from one or a few images [14, 6, 35, 15, 12], from layout information such as floor plans [58, 5], bird’s-eye views (semantic maps in top view) [28, 69], terrain maps [10] and 3D proxies [48], and as well as from text prompts [63, 49, 26, 54, 17]. However, each method has its own advantages and disadvantages in terms of scene control characteristics, and it is difficult to generate a 3D scene that appropriately reflects the intentions. We propose a method to address these challenges by integrating partial images, layout information, and text prompts as input conditions in a complementary manner. Furthermore, layout conditions were designed for each domain. Furthermore, while layout conditions need to be designed for each domain, the proposed method switches between converters for layout representations, enabling the generation of a variety of scenes from indoor to outdoor.
@@ 2.3 Monocular Depth Estimation
Monocular depth estimation involves estimating the depth of each pixel in a single RGB image. In recent years, deep learning-based methods have progressed significantly, and methods based on convolutional neural networks [47, 34, 32, 66, 67, 70, 38] and transformers [7, 13, 68, 55, 42] have been proposed. Monocular depth estimation for 360-degree images was also investigated [73, 33, 16, 59, 74, 60, 40, 43, 1]. In this study, we incorporated monocular depth estimation into a 360-degree RGB-D generation pipeline.
@@ 3.1.3 ERP Conversion
Figure 4: The pipeline of generating 360-degree RGB from a partial image, coarse depth map, semantic map, and text prompts. The partial image, coarse depth, and semantic map are embedded in the common latent space using encoders, and their channels are concatenated. Along with the text prompt, 360-degree RGB is generated in the framework of ControlNet[71] based on LDM[45].
@@
Fine-tuning of the base model degrades image-to-text performance. To mitigate this phenomenon, we additionally use a dataset with text annotations only for fine-tuning. If one model is trained for different combinations of conditions, the learning may not be generalized to other combinations of conditions. We introduce condition dropout (CD), in which training is performed by randomly changing the combination of conditions. Each condition is dropped with a probability of 50%, with the ERP image conditions being replaced by pixel values of 0 and text replaced by an empty string.
@@ 3.3.1 End-to-End Estimation
In the end-to-end approach, the depth is estimated using U-Net [46] with a self-attention mechanism [57] with four channels of RGB-D as the input, and one channel of depth as the output. The network is trained to minimize the L1 loss between the network outputs and ground truth. Details of the network configuration are provided in appendix A.
@@ 3.3.2 Depth Integration
where,$D_{k}=\tilde{d}_{k}^{\top}\{\Phi_{k}^{-1}+(\sum_{n=0\backslash k}^{N}\Phi_{n})^{-1}\}^{-1}\tilde{d}_{k},U_{k,l}=-\tilde{d}_{k}^{\top}\Phi_{k}(\sum_{n=0}^{N}\Phi_{n})^{-1}\Phi_{l}\tilde{d}_{l},b_{k}=\tilde{d}_{k}^{\top}\Phi_{k}(\sum_{n=0}^{N}\Phi_{n})^{-1}\Phi_{0}d_{0}$. The derivation of the equation and setting of weights $\{\Phi_{n}\}_{n=0}^{N}$ are described in appendix A.
@@ 4 Dataset
We build fine-tuning datasets for indoor and outdoor scenes, respectively. We create artificial datasets with layout annotations using computer graphics as the base dataset, whereas datasets without layout annotations are created using actual captured datasets as the auxiliary dataset.
@@ 4.1 Indoor Scene
For the auxiliary dataset, we used the Matterport 3D dataset [9], which is an indoor real-world 360 dataset captured by Matterport’s Pro 3D camera including 10,800 RGB-D panoramic images. Similar to the structured 3D dataset, partial images and text were annotated. The depth and semantic maps included in the dataset were not used, and zero was assigned as the default value for the coarse depth and semantic map during training. The data were divided into 7675 samples for training and 2174 samples for testing.
@@ 5 Experimental Results
Quantitative and qualitative experiments were conducted to verify the effectiveness of the proposed method for generating 3D scenes under multiple conditions.
@@ 5.1 Implementation Details
First, we evaluate 360-degree RGB generation. We used the peak-signal-to-noise-ratio (PSNR) as the evaluation metric: PSNR (whole) for the entire image between the ground truth and generated images, PSNR (parial) for the region of the partial image given by the input. We also emply the FID [24], which is a measure of the divergence of feature distributions between the ground truth and generated images, and the CLIP score (CS) [41, 23], which promptly quantifies the similarity with the input text. PSNR (whole) measures the reproducibility of the ground truth at the pixel value level, PSNR (partial) measures the degree to which the partial image given as the input is reflected in the generated result, FID measures the plausibility of the generated image, and CS measures the fidelity of the generated result to text prompts. Because there is no comparison method that uses partial images, layouts, and text prompts as inputs to generate a 360-degree image, we compared our method with PanoDiff [62], which is a state-of-the-art 360-degree RGB image generation model that uses partial images and texts. We implemented it and used PanoDiff with the encoder of the layout information removed in the proposed method for a fair comparison using the same network configurations and pretrained models.
@@
Ours (w/o CD) 12.56 35.39 18.87 30.72 12.46 34.68 29.54 29.71
Ours (w/ CD) 12.42 33.29 18.84 30.71 13.29 34.81 29.05 29.93
table 1 shows the quantitative evaluation results of 360-degree RGB generation on the Structured 3D dataset and the SceneDreamer dataset. The results of the proposed method are shown with and without CD. PanoDiff is superior in terms of PSNR (partial) and CS, which is a reasonable result since PanoDiff is a method that takes only partial images and text prompts as conditions for image generation. However, the proposed method is superior to PSNR (whole) and FID, which indicates that the reproducibility and plausibility of the generated images can be enhanced by considering layout information as a condition as well. Comparing the proposed method with and without CD, FID tended to be slightly better when CD was present, whereas PSNR (whole), PSNR (partial), and CS were superior or inferior depending on the two datasets. The better performance of CD on the SceneDreamer dataset can be attributed to the larger number of samples in the auxiliary dataset.
@@
fig. 7 shows the examples of generating a 360-degree RGB image for the test set of the Structured 3D dataset and the SceneDreamer dataset. PanoDiff, which does not use the layout information as a condition, generates images that differ significantly from the ground truth. This may have led to the degradation of PSNR (whole) and FID. Although the image generated by the proposed method differs from the ground-truth image at the pixel level, it can generate images with room geometry, terrain, and object placement in accordance with the given conditions. When comparing methods with and without CD, it is difficult to determine the perceived superiority in most cases. A closer look at fig. 7 (d) shows that additional objects, such as trees, are more likely to appear in Ours (w/o CD) on the SceneDreamer dataset, which may be a factor in the PSNR degradation.
Table 2: CS evaluation results for base model forgetting
$\checkmark$ 29.48 24.75
$\checkmark$ $\checkmark$ 29.34 26.24
$\checkmark$ $\checkmark$ $\checkmark$ 30.23 29.26
@@
Next, we evaluate the depth of the generated 360-degree image. Because the estimated depth has scale and offset degrees of freedom, its value was determined to minimize the squared error with the ground-truth depth, similar to the method presented in [42]. We used the root mean squared error (RMSE) and mean absolute value of the relative error, ${\rm AbsRel}=\frac{1}{M}\sum_{i=1}^{M}\frac{|z_{i}-z_{i}^{*}|}{z_{i}^{*}}$, where $M$ is the number of pixels, $z_{i}$ is the estimated depth of the $i$th pixel, and $z_{i}^{*}$ is the ground-truth depth of the $i$th pixel. Pixels at infinity were excluded from evaluation. table 3 shows the results of the quantitative evaluation of depth generation on the Structured 3D dataset and the SceneDreamer dataset. For comparison, the results of 360MonoDepth which is a 360 monocular depth estimation [43] method; LeRes (ERP), which is LeRes [70] directly applied to ERP; and LeRes (multi views), which applies LeRes to multiple tangent images of a 360-degree image and integrates the estimated depths in a section 3.3 manner without using coarse depth, are also shown. In terms of RMSE and AbsRel, our method (end-to-end) was the best for the structured 3D dataset, and our method (depth integration) was the best for the SceneDreamer dataset. It was also shown that combining LeRes with coarse depth increased accuracy compared to using LeRes alone. The end-to-end method is relatively ineffective for the SceneDreamer dataset. This may be because the number of samples in the dataset was small and the depth was estimated to be close to the coarse depth.
@@ 5.4 Results in the Wild
Finally, we evaluated the results of 3D scene generation based on user-generated conditions outside the dataset used for fine-tuning. Examples of 3D scenes generated by the proposed method, conditioned on partial images, layouts, and text, are shown in figs. 1 and 9. These conditions were created freely by the authors. It can be seen that the generated scene contains the given partial image and conforms to the instructions of the text prompt according to the given layout. These results show that the proposed method can generate 3D scenes with the appearance, geometry, and overall context controlled according to the input information, even outside the dataset used for fine-tuning. Details of the experimental setup, additional samples, ablation studies, and limitations are described in appendices B and C.
@@ A.1 End-toEnd Network Configuration
The structure of the network that generates a fine depth from a coarse depth and the generated RGB end-to-end is shown in figs. 10 and 11. The network consists of a combination of U-Net [46] and self-attention [57], with four channels of RGB-D as the input and one channel of depth as the output. The network was trained to minimize the L1 loss between the depth output from the network and the depth of the ground truth. The model was trained from scratch using the Adam optimizer with a learning rate of $4.5\times 10^{-6}$ and a batch size of six.
@@ Appendix B Additional Results
B.1 360-Degree RGB Generation
@@
figs. 13 and 14 show additional samples of 360-degree RGB image generation for the Structured 3D dataset and SceneDreamer dataset, respectively.
B.2 Generation Results from Subset of Conditions
Conditions RGB Depth
[…]
$\checkmark$ $\checkmark$ 12.04 34.46 43.86 28.19 8.96 0.0100
$\checkmark$ $\checkmark$ 11.45 - 21.71 30.67 8.78 0.0056
$\checkmark$ $\checkmark$ 11.48 33.64 21.83 30.93 24.56 0.0172
$\checkmark$ 11.40 35.00 55.08 27.00 23.94 0.0158
$\checkmark$ 11.12 - 59.70 27.59 5.02 0.0086
$\checkmark$ 10.67 - 25.90 30.85 24.53 0.0171
10.43 - 87.69 24.40 24.00 0.0180
To verify the contribution and robustness of each condition of the proposed method, experiments were conducted to generate 360-degree RGB-D from a subset of partial images, layouts, and text prompts. Generation was performed using the proposed method with the CD for the test set of the structured 3D dataset. Because depth estimation in the proposed method requires layout information, LeRes (ERP) [70], a monocular depth estimation of ERP images, was used in the absence of layout conditions. table 4 shows the values of each evaluation metric for the generated results. In terms of FID, it can be seen that the proposed method does not significantly degrade performance when text conditions are included in the generation conditions. This is largely owing to the performance of the text-to-image model used as the base model to ensure the plausibility of the generated image. However, PSNR (whole) decreases in the absence of partial image and layout conditions, indicating that the contribution of these conditions to the composition of the overall structure is high. In addition, CS naturally decreases without the text condition. However, even without the text condition, CS is larger than that in the unconditional generation case, indicating that semantic reproduction is possible to some extent, even from partial images and layout information. For depth generation, the accuracy is significantly degraded because it is impossible to use depth estimation with a coarse depth in the absence of layout conditions. When generated from partial images and text, its performance was comparable to PanoDiff.
B.3 Results in the Wild
@@
Additional examples of 3D scenes generated using the proposed method conditioned on text, partial images, and layouts are presented in figs. 15, 16, 17, 18 and 19. In these figures, the aspect ratios of the ERP images were converted to 2:1 for display purposes. These conditions were created freely by the authors. It can be seen that the generated scene contains the given partial image and conforms to the instructions of the text prompt according to the given layout. In addition to the coarse depth created by the room shape or terrain alone, the geometry of objects such as chairs, tables, trees, and buildings can be seen. fig. 15 shows how various scenes can be generated in a controlled manner by changing the combination of layout and text for the same partial image. figs. 16, 17, 18 and 19 shows that our method can generate a variety of 3D scenes from photos on the web, photos taken in the real world, and fanciful paintings, taking into account the layout and text requirements we give. These results show that the proposed method can generate 360-degree RGB-D images with appearance, geometry, and overall context controlled according to the input information, even outside the dataset used for fine-tuning.

### v2

@@ 2 Related Work
3D scene generation involves the creation of a model of a 3D space that includes objects and backgrounds, based on user-specified conditions. In recent years, the use of generative models, such as VAE [31, 45], GAN [19], autoregressive models [58], and diffusion models [51, 26], has made rapid progress. There are methods to generate a 3D scene from random variables [38, 8], from one or a few images [14, 6, 36, 15, 12], from layout information such as floor plans [59, 5], bird’s-eye views (semantic maps in top view) [29, 70], terrain maps [10] and 3D proxies [49], and as well as from text prompts [64, 50, 27, 55, 17]. However, each method has its own advantages and disadvantages in terms of scene control characteristics, and it is difficult to generate a 3D scene that appropriately reflects the intentions. We propose a method to address these challenges by integrating partial images, layout information, and text prompts as input conditions in a complementary manner. Furthermore, while layout conditions need to be designed for each domain, the proposed method switches between converters for layout representations, enabling the generation of a variety of scenes from indoor to outdoor.
@@ 2.3 Monocular Depth Estimation
Monocular depth estimation involves estimating the depth of each pixel in a single RGB image. In recent years, deep learning-based methods have progressed significantly, and methods based on convolutional neural networks [48, 35, 33, 67, 68, 71, 39] and transformers [7, 13, 69, 56, 43] have been proposed. Monocular depth estimation for 360-degree images was also investigated [74, 34, 16, 60, 75, 61, 41, 44, 1]. However, since the accuracy of conventional monocular depth estimation is not sufficient, this study aims to improve accuracy by combining it with layout conditions.
@@ 3.1.3 ERP Conversion
Figure 4: The pipeline of generating 360-degree RGB from a partial image, coarse depth map, semantic map, and text prompts.
@@ 3.3.1 End-to-End Estimation
In the end-to-end approach, the depth is estimated using U-Net [47] with a self-attention mechanism [58] with four channels of RGB-D as the input, and one channel of depth as the output. The network is trained to minimize the L1 loss between the network outputs and ground truth. Details of the network configuration are provided in section A.1.
@@ 3.3.2 Depth Integration
where, $D_{k}$ $=$ $\tilde{d}_{k}^{\top}\{\Phi_{k}^{-1}+(\sum_{n=0\backslash k}^{N}\Phi_{n})^{-1}\}^{-1}\tilde{d}_{k}$, $U_{k,l}$ $=$ $-\tilde{d}_{k}^{\top}\Phi_{k}(\sum_{n=0}^{N}\Phi_{n})^{-1}\Phi_{l}\tilde{d}_{l},b_{k}=\tilde{d}_{k}^{\top}\Phi_{k}(\sum_{n=0}^{N}\Phi_{n})^{-1}\Phi_{0}d_{0}$. The derivation of the equation and setting of weights $\{\Phi_{n}\}_{n=0}^{N}$ are described in section A.2.
@@ 4 Dataset
We fine-tune our model using the following two types of datasets for indoor and outdoor scenes, respectively. We create artificial datasets with layout annotations using computer graphics as the base dataset, whereas datasets without layout annotations are created using actual captured datasets as the auxiliary dataset.
@@ 4.1 Indoor Scene
For the auxiliary dataset, we used the Matterport 3D dataset [9], which is an indoor real-world 360 dataset including 10,800 RGB-D panoramic images. Similar to the structured 3D dataset, partial images and text were annotated. The depth and semantic maps included in the dataset were not used, and zero was assigned as the default value for the coarse depth and semantic map during training. The data were divided into 7675 samples for training and 2174 samples for testing.
@@ 5 Experimental Results
Quantitative and qualitative experiments were conducted to verify the effectiveness of the proposed method, MaGRITTe, for generating 3D scenes under multiple conditions.
@@ 5.1 Implementation Details
First, we evaluate 360-degree RGB generation. Because there is no comparison method that uses partial images, layouts, and text prompts as inputs to generate a 360-degree image, we compared our method with PanoDiff [63], which is a state-of-the-art 360-degree RGB image generation model that uses partial images and texts. We implemented it and used PanoDiff with the encoder of the layout information removed in MaGRITTe for a fair comparison using the same network configurations and pretrained models.
@@
MaGRITTe (ours) 12.56 35.39 18.87 30.72 13.29 34.81 29.05 29.93
Table 2: Evaluation results for object type and placement. Note that the object positions in the input condition are given by the bounding boxes as shown in fig. 5, therefore even in ground truth images, it doesn’t match perfectly.
Method Precision Recall IoU
Ground truth 0.482 0.349 0.284
PanoDiff 0.245 0.170 0.124
MaGRITTe (ours) 0.424 0.273 0.227
table 1 shows the quantitative evaluation results of 360-degree RGB generation on the Structured 3D dataset and the SceneDreamer dataset. We used the peak-signal-to-noise-ratio (PSNR) as the evaluation metric: PSNR (whole) for the entire image between the ground truth and generated images, PSNR (parial) for the region of the partial image given by the input. We also emply the FID [25], which is a measure of the divergence of feature distributions between the ground truth and generated images, and the CLIP score (CS) [42, 24], which promptly quantifies the similarity with the input text. PanoDiff is superior in terms of PSNR (partial) and CS, which is a reasonable result since PanoDiff is a method that takes only partial images and text prompts as conditions for image generation. However, MaGRITTe is superior to PSNR (whole) and FID, which indicates that the reproducibility and plausibility of the generated images can be enhanced by considering layout information as a condition as well.
table 2 shows the results of the evaluation of the controllability of object type and placement. Semantic segmentation [20] was performed on the 360-degree images generated for Structured3D dataset to evaluate precision, recall, and IoU for bounding boxes in the input conditions. MaGRITTe is superior to PanoDiff and produces results closer to the ground truth images, indicating that the condition-aware object placement is realized.
@@
fig. 7 shows the examples of generating a 360-degree RGB image for the test set of the Structured 3D dataset and the SceneDreamer dataset. PanoDiff, which does not use the layout information as a condition, generates images that differ significantly from the ground truth. This may have led to the degradation of PSNR (whole) and FID. Although the image generated by MaGRITTe differs from the ground-truth image at the pixel level, it can generate images with room geometry, terrain, and object placement in accordance with the given conditions.
@@
Ours (w/o coarse depth) 9.837 0.0070 15.28 0.0196
Next, we evaluate the depth of the generated 360-degree image. Because the estimated depth has scale and offset degrees of freedom, its value was determined to minimize the squared error with the ground-truth depth, similar to the method presented in [43]. We used the root mean squared error (RMSE) and mean absolute value of the relative error, ${\rm AbsRel}=\frac{1}{M}\sum_{i=1}^{M}\frac{|z_{i}-z_{i}^{*}|}{z_{i}^{*}}$, where $M$ is the number of pixels, $z_{i}$ is the estimated depth of the $i$th pixel, and $z_{i}^{*}$ is the ground-truth depth of the $i$th pixel. Pixels at infinity were excluded from evaluation. table 3 shows the results of the quantitative evaluation of depth generation on the Structured 3D dataset and the SceneDreamer dataset. For comparison, the results of 360MonoDepth which is a 360 monocular depth estimation [44] method; LeRes (ERP), which is LeRes [71] directly applied to ERP; and LeRes (multi views), which applies LeRes to multiple tangent images of a 360-degree image and integrates the estimated depths in a section 3.3 manner without using coarse depth, are also shown. In terms of RMSE and AbsRel, our method (end-to-end) was the best for the structured 3D dataset, and our method (depth integration) was the best for the SceneDreamer dataset. It was also shown that combining LeRes with coarse depth increased accuracy compared to using LeRes alone. Ours (w/o coarse depth) is an end-to-end depth estimation method that uses only RGB without the coarse depth, and we can see that the accuracy is lower than when using coarse depth in the Structured3D dataset. The end-to-end method is relatively ineffective for the SceneDreamer dataset. This may be because the number of samples in the dataset was small and the depth was estimated to be close to the coarse depth.
@@ 5.4 Results in the Wild
We evaluated the results of 3D scene generation based on user-generated conditions outside the dataset used for fine-tuning. Examples of 3D scenes generated by MaGRITTe, conditioned on partial images, layouts, and text, are shown in figs. 1 and 8. These conditions were created freely by the authors. It can be seen that the generated scene contains the given partial image and conforms to the instructions of the text prompt according to the given layout. These results show that MaGRITTe can generate 3D scenes with the appearance, geometry, and overall context controlled according to the input information, even outside the dataset used for fine-tuning.
5.5 Generation Results from Subset of Conditions
Conditions RGB Depth
[…]
$\checkmark$ $\checkmark$ 12.04 34.46 43.86 28.19 8.96 0.0100
$\checkmark$ $\checkmark$ 11.45 - 21.71 30.67 8.78 0.0056
$\checkmark$ $\checkmark$ 11.48 33.64 21.83 30.93 24.56 0.0172
$\checkmark$ 11.40 35.00 55.08 27.00 23.94 0.0158
$\checkmark$ 11.12 - 59.70 27.59 5.02 0.0086
$\checkmark$ 10.67 - 25.90 30.85 24.53 0.0171
10.43 - 87.69 24.40 24.00 0.0180
To verify the contribution and robustness of each condition of the proposed method, experiments were conducted to generate 360-degree RGB-D from a subset of partial images, layouts, and text prompts. Generation was performed for the test set of the structured 3D dataset. Because depth estimation in MaGRITTe requires layout information, LeRes (ERP) [71], a monocular depth estimation of ERP images, was used in the absence of layout conditions. table 4 shows the values of each evaluation metric for the generated results. In terms of FID, it can be seen that MaGRITTe does not significantly degrade performance when text conditions are included in the generation conditions. This is largely owing to the performance of the text-to-image model used as the base model to ensure the plausibility of the generated image. However, PSNR (whole) decreases in the absence of partial image and layout conditions, indicating that the contribution of these conditions to the composition of the overall structure is high. In addition, CS naturally decreases without the text condition. However, even without the text condition, CS is larger than that in the unconditional generation case, indicating that semantic reproduction is possible to some extent, even from partial images and layout information. For depth generation, the accuracy is significantly degraded because it is impossible to use depth estimation with a coarse depth in the absence of layout conditions. When generated from partial images and text, its performance was comparable to PanoDiff. Details of the experimental setup, additional samples, ablation studies, and limitations are described in appendices B and C.
@@ A.1 End-toEnd Network Configuration
The structure of the network that generates a fine depth from a coarse depth and the generated RGB end-to-end is shown in figs. 9 and 10. The network consists of a combination of U-Net [47] and self-attention [58], with four channels of RGB-D as the input and one channel of depth as the output. The network was trained to minimize the L1 loss between the depth output from the network and the depth of the ground truth. The model was trained from scratch using the Adam optimizer with a learning rate of $4.5\times 10^{-6}$ and a batch size of six.
@@ Appendix B Additional Results
B.1 Condition Dropout
Fine-tuning of the base model degrades image-to-text performance. To mitigate this phenomenon, we additionally use the auxiliary dataset (see Section 4) with text annotations only for fine-tuning. If one model is trained for different combinations of conditions, the learning may not be generalized to other combinations of conditions. We introduce condition dropout (CD), in which training is performed by randomly changing the combination of conditions. Each condition is dropped with a probability of 50%, with the ERP image conditions being replaced by pixel values of 0 and text replaced by an empty string.
table 5 shows the results of comparing the presence or absence of CD in the proposed method. FID tended to be slightly better when CD was present, whereas PSNR (whole), PSNR (partial), and CS were superior or inferior depending on the two datasets. The better performance of CD on the SceneDreamer dataset can be attributed to the larger number of samples in the auxiliary dataset.
Structured3D dataset SceneDreamer dataset
[…]
w/o CD 12.56 35.39 18.87 30.72 12.46 34.68 29.54 29.71
w/ CD 12.42 33.29 18.84 30.71 13.29 34.81 29.05 29.93
Table 6: CS evaluation results for base model forgetting
$\checkmark$ 29.48 24.75
$\checkmark$ $\checkmark$ 29.34 26.24
$\checkmark$ $\checkmark$ $\checkmark$ 30.23 29.26
B.2 Comparison with Text2Room
Text2Room [27] is a method for generating 3D scenes as meshes by repeatedly generating images in multiple viewpoints from the input text. This method can also be used to control the layout of the generated 3D scene by changing the input text according to the viewpoint. However, layout guided generation in Text2Room is different from our setting, because it changes the text prompts for the direction of observation and cannot take geometric shapes as conditions. fig. 13 shows an example of a scene generated by Text2Room under the same conditions as in fig. 1. Text2Room is less accurate in the placement of objects and is unable to generate room shapes to suit the conditions. Conditioning the layout with semantic map and coarse depth is the advantage of our method.
Figure 13: Comparison with Text2Room. (a) ERP images of the generated 3D scenes, (b) Room shapes in the top view.
B.3 360-Degree RGB Generation
@@ B.2 Comparison with Text2Room
figs. 14 and 15 show additional samples of 360-degree RGB image generation for the Structured 3D dataset and SceneDreamer dataset, respectively.
B.4 Results in the Wild
@@
Additional examples of 3D scenes generated using the proposed method conditioned on text, partial images, and layouts are presented in figs. 16, 17, 18, 19 and 20. In these figures, the aspect ratios of the ERP images were converted to 2:1 for display purposes. These conditions were created freely by the authors. It can be seen that the generated scene contains the given partial image and conforms to the instructions of the text prompt according to the given layout. In addition to the coarse depth created by the room shape or terrain alone, the geometry of objects such as chairs, tables, trees, and buildings can be seen. fig. 16 shows how various scenes can be generated in a controlled manner by changing the combination of layout and text for the same partial image. figs. 17, 18, 19 and 20 shows that our method can generate a variety of 3D scenes from photos on the web, photos taken in the real world, and fanciful paintings, taking into account the layout and text requirements we give. These results show that the proposed method can generate 360-degree RGB-D images with appearance, geometry, and overall context controlled according to the input information, even outside the dataset used for fine-tuning.